Grade 10Physical Chemistry

Henry’s law constant for CO2 in water is 1.67 X 108 Pa at 298 K.
Calculate the quantity of CO2 in 500 ml of soda water when packed under 2.5 atm CO2 pressure at 298 K.

Aiden Mathew Denny
7 Years agoGrade 10

1 Answer

Arun
ApprovedApproved Tutor Answer7 Years ago
According to Henry's law, pCO2 = KHx
Here x is the mole fraction of carbon dioxide in the carbon dioxide water mixture.
pCO2 = 2.5 atm = 2.5 * 1.01325 * 105 Pa = 2.533125 * 105 Pa
x = mole fraction = 2.533125 * 105/ 1.67 * 108 = 0.00152
x = moles of CO2/(moles of CO2 + moles of H2O) ≈moles of CO2/moles of H2Oas mole of CO2are negligible in comparison to the number of moles of water
500 ml of water = 500g of water = 500/18 = 27.78 moles of water.
So, moles of CO2/27.78 = 0.00152 or moles of CO2 = 0.042 moles.
 
Hence the amount of CO2 in 500 ml of soda water is 0.042 * 44 = 1.848 g
 
Regards
Arun (askIITians forum expert)